OpenStudy (anonymous):

Cal 2 int. By Parts question?

OpenStudy (anonymous):

\[\int\limits x ^{12}Ln(x ^{7})dx\]

OpenStudy (anonymous):

can you show me how to do the steps?

OpenStudy (saifoo.khan):

Try it out, Paulie!

OpenStudy (anonymous):

\[\int\limits x ^{12+1} \frac{ 1 }{ 12+1? }=\frac{ 1 }{ 13}x ^{13}Ln(x ^{7})dx\]

OpenStudy (anonymous):

Am I on the right track for the first part?

OpenStudy (saifoo.khan):

@Directrix

OpenStudy (saifoo.khan):

Sorry, i gtg.

OpenStudy (anonymous):

here is step-by-step on how i attempted this problem :O